OpenStudy (anonymous):

Evaluate: 7 z +negative z , when z = 4. A. 24 B. 0 C. -112 D. -7

jigglypuff314 (jigglypuff314):

when you add a negative you subtract so 7z + (-z) = 7z - z = 6z then plug in 4 for z so 6z = 6(4) = 6 times 4 = ?

OpenStudy (anonymous):

use a calculator :P

jigglypuff314 (jigglypuff314):

yeah, just copy paste XD http://www.wolframalpha.com/input/?i=7+z+%2Bnegative+z+%2C+when+z+%3D+4

OpenStudy (anonymous):

6 * 4 = ? 24